# Dynamical Tidal Response of Schwarzschild Black Holes

### v1

@@ 3.2.2 Even sector
$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}+\langle E_{+\,rr}\rangle_{\text{in-in}}$ $=-c_{\text{ext}}\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\ell(\ell-1)Y_{\ell}^{m}$ (3.103)
$\!\!\!\times\left(r^{\ell-2}+\mu^{2\varepsilon}K^{(E)}_{\ell}(\omega)\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+1}\right),$
@@
$E_{rr}=-\partial_{r}^{2}H_{0}-2i\omega\partial_{r}H_{1}.$ (3.106)
@@
$\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\ell!\,\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}K^{(E)}_{\ell}(\omega)=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.109)
@@
$\frac{45}{2\pi}K^{(E)}_{\ell=2}(\omega)=i\omega r_{s}^{6}-\omega^{2}r_{s}^{7}\left(-\frac{797}{1260}+\log(\mu r_{s})\right).$ (3.118)
Repeating the same procedure for $\ell=3$ and $\ell=4$ (see Appendix C for details) we get:
$\frac{18900}{2\pi}K^{(E)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}-\omega^{2}r_{s}^{9}\left(-\frac{709}{840}+\log(\mu r_{s})\right),$ (3.119) $\frac{11113200}{2\pi}K^{(E)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}-\omega^{2}r_{s}^{11}\left(-\frac{5\,501}{5\,775}+\log(\mu r_{s})\right).$
@@ 3.3 Symmetries
Aside from the fact that they are nonzero, the dynamical Love numbers that we have computed have some properties worth remarking on. Most notably, if we expand the response kernels $K$ as in (3.53) and (3.94) and then use (3.87) and (3.118) along with (3.88) and (3.119) to determine the effective worldline couplings, we see that $\lambda_{1,\ell}^{E}=\lambda_{1,\ell}^{B}$ and the coefficients of the logs in $\lambda_{2,\ell}^{E}$ are exactly the same as those in $\lambda_{2,\ell}^{B}$. [^28] In making this comparison, there is a relative normalization factor of $1/2$ in the magnetic $\lambda$ coefficients coming from the relation between the Weyl tensor and gravito-magnetic field. That is, the dynamical Love numbers in the even and odd sectors are equal—this can be understood as a consequence of Chandrasekhar’s duality.
As we mentioned in Section 3.1.2, Chandrasekhar discovered a mapping between the Regge–Wheeler and Zerilli equations [76, 77, 78]. There we used this fact as a solution-generating technique to obtain solutions in the Zerilli sector from those in the Regge–Wheeler sector. Here we note that this relation can also be read as a symmetry [7]. Specifically, both the Regge–Wheeler potential and Zerilli potential in $D=4$ can be written as
@@
This is an off-shell symmetry with a corresponding Noether current. At large distances it interchanges the electric and magnetic parts of the Weyl tensor, and so is a Schwarzschild generalization of electric-magnetic duality. [^29] Note that it keeps the background fixed, it does not map Schwarzschild into Taub–NUT. It is therefore directly a symmetry of linearized perturbations around Schwarzschild.
In the present context, this symmetry maps the worldline response operators (3.40) in the electric and magnetic sectors into each other, relating their coefficients. The astute reader, however, will note that the finite contributions in e.g., (3.87) and (3.118) differ. This would seem to be in conflict with the symmetry. However, we computed these quantities via dimensional regularization, and the symmetry (3.121) exists only in $D=4$. Since we have chosen a regularization scheme that breaks the symmetry, it is not surprising that the final answers do not respect it. Instead, the existence of the symmetry suggests that a scheme exists that preserves the symmetry. Indeed, an important consistency check is that the scheme-independent parts of the response—the imaginary dissipative response and the coefficient of the logarithms—match exactly between the two sectors. In addition, we can certainly pick the scales $\mu$ in the two sectors differently so that the finite pieces agree, which would be a scheme that preserves the duality.
@@ 4 Conclusions
What we find after matching to define the dynamical Love numbers of black holes is that they have several interesting properties. They are nonzero, and have a characteristic running, where the coefficient of the logarithm in all cases is $\omega r_{s}$ times the corresponding dissipative response coefficient. One way to read this relation is that the beta function for the dynamical Love numbers is determined by the dissipative response. However, this appears to become more complicated at subleading order [74]. Another way to read this relation is that the discontinuity of the conservative part of $K$ is $\beta\omega/2$ times the dissipative part of $K$, where $\beta=4\pi r_{s}$ is the inverse black hole temperature. [^30] This looks suspiciously like a fluctuation–dissipation relation. However, it is a relation involving just the causal Green’s function. (Interestingly this relation continues to hold for the most transcendental parts of the response kernel at subleading orders in frequency, at least in the examples of [74].) It would be nice to understand the features of these logarithmic terms better. Another interesting feature is that, as expected from the symmetries of black hole perturbations, there exists a scheme in which the electric and magnetic dynamical Love numbers (along with the dissipative responses) are equal.
There are a number of future directions suggested by this study that it would be interesting to pursue. Perhaps the most obvious would be to extend the computation done here to general $\ell$ and (relatedly) to higher dimension. In principle there are only technical complications to doing so, and it would be very interesting to see if the patterns seen in the four-dimensional case survive there. Relatedly, one could carry out the computation of dynamical responses to external electromagnetic fields and consider the generalization to charged black hole responses. Perhaps the most phenomenologically interesting generalization would be to compute the full dynamical Love numbers for a spinning Kerr black hole. To do this, the main advances required are technical. In particular, re-summing and regulating gravitational effects in the EFT is most simply done in dimensional regularization, but gravitational couplings to spinning black holes in higher dimensions are somewhat complicated. At leading order, we can extract a Kerr black hole’s dynamical Love number from those of a Schwarzschild black hole by exploiting the fact that a Kerr black hole is a spinning Schwarzschild black hole. The coordinate transformation between the (rotating) black hole’s rest frame and the laboratory frame captures the leading-in-spin effect (much in the same way that a Kerr black hole’s zero frequency dissipation is related to the finite frequency Schwarzschild dissipation) [46, 19]. Nevertheless the computation of the full response for Kerr is a problem that we hope to return to.
@@ Acknowledgements:
Thanks to John Joseph Carrasco, Lam Hui, Riccardo Penco, and Mikhail Solon for helpful conversations AJ and DG are supported in part by DOE (HEP) Award DE-SC0025323. The work of MJR is supported through the NSF grant PHY-2309270. The research of LS has been funded by the French National Research Agency (ANR) under project ANR-24-CE31-1097-01. The work is supported by a PhD Joint Program of the International Research Center for Fundamental Scientific Discovery (IRC Discovery).
@@ Appendix A Scalar dynamical response running
with $m$ a non-negative integer. [^31] In (A.8), the lower bound in the integral (which we chose to be $1$ for convenience) is completely immaterial, as this simply amounts to a redefinition of the integration constants in (2.21). Note that no $\log(x)/x^{\ell+1}$ scaling is contained in the homogeneous solutions of (2.17). From the properties of the hypergeometric function, each of these terms is of the form $\text{pol}_{1}(x)+\text{pol}_{2}(x)\times\log(x)$, at large $x$, where $\text{pol}_{1}(x)$ and $\text{pol}_{2}(x)$ are generic ordinary polynomials with positive powers of $x$. In conclusion, no $\log(x)/x^{\ell+1}$ term results from the $x\rightarrow\infty$ expansion of the $I_{n}$ integrals (A.6), for all $n$.
The same conclusion is trivially true for the second integral in (A.4), since both $P_{\ell}(y)$ and $W_{\ell-1}(y)$ are polynomials in (y1)(y-1). Therefore, there is no $\log(x)/x^{\ell+1}$ type of term resulting from the first integral in (2.21). Let us then focus on the second integral in (2.21):
@@
$\frac{B}{4}\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)^{2}=2B\log(x-1)+\text{pol}(x).$ (A.10)
@@ B.1 Far-zone scalar solution
$_{p}\tilde{F}_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right]\equiv\frac{\,{}_{p}F_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right]}{\Gamma(b_{1})\cdots\Gamma(b_{q})}.$ (B.17)
@@
To expand the Coulomb functions, we can then use their relations to spherical Bessel functions, i.e.,
@@
Plugging this into the definition of $\phi^{(1)}_{\ell,\text{FZ}}=\frac{\omega}{z}(\xi^{(1)}_{\ell}+\frac{1}{2z}\xi^{(0)}_{\ell})$, we obtain the following $z\rightarrow 0$ behavior for $\phi_{\ell,\text{FZ}}^{(1)}$ [^36] We use $y_{0}^{\ell}=(2\ell-1)!!=\frac{(2\ell+1)!!}{2\ell+1}$ to simplify the expressions.
@@
$\quad+c_{1}^{(1)}j_{0}^{\ell}z^{\ell}+O(z^{\ell+1})\Bigg]$
$=$ $\,\,\omega c_{2}^{(1)}\frac{(2\ell+1)!!}{2\ell+1}z^{-\ell-1}\left(1+\sum_{n=1}^{\ell-1}\frac{y_{n}^{\ell}}{y_{0}^{\ell}}z^{2n}\right)-\omega c_{1}^{(0)}\frac{\ell}{2(2\ell+1)!!}z^{\ell-1}$
$+\frac{\omega c_{1}^{(1)}}{(2\ell+1)!!}z^{\ell}+O(z^{\ell+1}).$ (B.27)
@@ C.1 General relativistic solutions
$\penalty 10000\ \hskip 78.24507pt+\frac{13\,923}{500\,000}\frac{r_{s}^{2}}{r^{2}}+\frac{1\,622\,851}{105\,000\,000}\frac{r_{s}^{3}}{r^{3}}\bigg)$
$\penalty 10000\ \hskip 28.45274pt+C\omega^{2}r_{s}^{2}\bigg[-\frac{1}{3}\frac{r^{6}}{r_{s}^{6}}-\frac{11}{10}\frac{r^{5}}{r_{s}^{5}}-\frac{r^{4}}{r_{s}^{4}}\!\left(\pi^{2}+\frac{51\,917}{2\,100}-\frac{13}{7}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r^{3}}{r_{s}^{3}}\!\left(\frac{19\pi^{2}}{30}+\frac{399\,451}{21\,000}-\frac{247}{210}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r^{2}}{r_{s}^{2}}\!\left(\frac{31\pi^{2}}{100}+\frac{1\,634\,147}{210\,000}-\frac{403}{700}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt-\frac{r}{10^{3}r_{s}}\!\left(93\pi^{2}+\frac{2\,039\,147}{700}-\frac{1\,209}{7}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt-\frac{221\pi^{2}}{10\,000}-\frac{10\,422\,677}{21\,000\,000}+\frac{2\,873}{70\,000}\log\frac{r_{s}}{r}$
$\penalty 10000\ \hskip 78.24507pt+\frac{r_{s}}{10^{5}r}\!\left(663\pi^{2}+\frac{21\,972\,677}{700}-\frac{8\,619}{7}\log\!\left(\frac{r_{s}}{r}\right)\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r_{s}^{2}}{10^{6}r^{2}}\!\left(-1\,989\pi^{2}+\frac{220\,495\,907}{2\,100}+\frac{25\,857}{7}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r_{s}^{3}}{10^{7}r^{3}}\!\left(5\,967\pi^{2}+\frac{58\,238\,757\,859}{44\,100}-\frac{5\,232\,713}{21}\log\frac{r_{s}}{r}\right)\bigg]+O\!\left(\frac{r_{s}^{4}}{r^{4}}\right),$
@@
$\penalty 10000\ \hskip 28.45274pt+Ci\omega r_{s}\bigg[\frac{1\,246}{15}\frac{r^{5}}{r_{s}^{5}}-\frac{10\,591}{90}\frac{r^{4}}{r_{s}^{4}}+\frac{15\,931}{540}\frac{r^{3}}{r_{s}^{3}}+\frac{32\,129}{3\,240}\frac{r^{2}}{r_{s}^{2}}-\frac{32\,129}{19\,440}\frac{r}{r_{s}}$
[…]
$\penalty 10000\ \hskip 28.45274pt+C\omega^{2}r_{s}^{2}\bigg[-\frac{14}{11}\frac{r^{7}}{r_{s}^{7}}-\frac{931}{330}\frac{r^{6}}{r_{s}^{6}}-\frac{r^{5}}{r_{s}^{5}}\!\left(\frac{14\pi^{2}}{3}+\frac{2\,798\,434}{17\,325}-\frac{3\,142}{495}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt+\frac{r^{4}}{r_{s}^{4}}\!\left(\frac{119\pi^{2}}{18}+\frac{202\,591\,567}{831\,600}-\frac{26\,707}{2\,970}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 78.24507pt-\frac{r^{3}}{r_{s}^{3}}\!\left(\frac{179\pi^{2}}{108}+\frac{332\,182\,607}{4\,989\,600}-\frac{281\,209}{124\,740}\log\frac{r_{s}}{r})\right)$
$\penalty 10000\ \hskip 78.24507pt-\frac{r^{2}}{r_{s}^{2}}\!\left(\frac{361\pi^{2}}{648}+\frac{603\,088\,753}{29\,937\,600}-\frac{567\,131}{748\,440}\log\frac{r_{s}}{r}\right)$

### v2

@@ 3.2.2 Even sector
$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}+\langle E_{+\,rr}\rangle_{\text{in-in}}$ $=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}c_{\rm ext}\ell(\ell-1)Y_{\ell}^{m}$ (3.103)
$\!\!\!\times\left(r^{\ell-2}+\mu^{2\varepsilon}\frac{K^{(E)}_{\ell}(\omega)}{M_{\rm Pl}^{2}}\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+1}\right),$
@@
$E_{rr}=-\frac{1}{M_{\rm Pl}}\left(\partial_{r}^{2}H_{0}+2i\omega\partial_{r}H_{1}\right).$ (3.106)
@@
$\frac{D-3}{D-2}\frac{(\ell+D-3)(\ell+D-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\ell!\,\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{K^{(E)}_{\ell}(\omega)}{M_{\rm Pl}^{2}}=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.109)
@@
$\frac{45}{2\pi M_{\rm Pl}^{2}}K^{(E)}_{\ell=2}(\omega)=i\omega r_{s}^{6}-\omega^{2}r_{s}^{7}\left(-\frac{797}{1260}+\log(\mu r_{s})\right).$ (3.118)
Note that the linear term in the frequency and the coefficient of the logarithm agree with [49] (see also [40]) after accounting for the different conventions. [^33] In particular, in the notation of [49], $\tilde{F}(\omega)=4K^{(E)}_{\ell=2}(\omega)$, where the factor of $4$ follows from canonical normalization of the graviton field. Repeating the same procedure for $\ell=3$ and $\ell=4$ (see Appendix C for details) we get:
$\frac{18900}{2\pi M_{\rm Pl}^{2}}K^{(E)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}-\omega^{2}r_{s}^{9}\left(-\frac{709}{840}+\log(\mu r_{s})\right),$ (3.119) $\frac{11113200}{2\pi{M_{\rm Pl}^{2}}}K^{(E)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}-\omega^{2}r_{s}^{11}\left(-\frac{5\,501}{5\,775}+\log(\mu r_{s})\right).$
@@ 3.3 Symmetries
Aside from the fact that they are nonzero, the dynamical Love numbers have some properties worth remarking on. Most notably, if we expand the response kernels $K$ as in (3.53) and (3.94) and then use (3.87) and (3.118) along with (3.88) and (3.119) to determine the effective worldline couplings, we see that $\lambda_{1,\ell}^{E}=\lambda_{1,\ell}^{B}$ and the coefficients of the logs in $\lambda_{2,\ell}^{E}$ are exactly the same as those in $\lambda_{2,\ell}^{B}$. [^34] In making this comparison, there is a relative normalization factor of $\ell/(\ell+1)$ in the definition of the magnetic $\lambda$ coefficients coming from the relation between the Weyl tensor and gravito-magnetic field. In addition, recall that in flat space Chandrasekhar’s duality acts by rotating the Riemann tensor and its dual, so that in particular $\delta E_{ab}=-\frac{1}{2}{\epsilon_{0a}}^{cd}B_{bcd}$ and $\delta B_{abc}={\epsilon^{0d}}_{bc}E_{ad}$ (see e.g. (3.72) and (3.103)). That is, the dynamical Love numbers in the even and odd sectors are equal. This can be understood as a consequence of Chandrasekhar’s duality.
As we mentioned in Section 3.1.2, Chandrasekhar discovered a mapping between the Regge–Wheeler and Zerilli equations [76, 77, 78]. There we used this fact as a solution-generating technique to obtain solutions in the Zerilli sector from those in the Regge–Wheeler sector. Here we note that this relation can also be read as a symmetry [7, 126]. Specifically, both the Regge–Wheeler potential and Zerilli potential in $D=4$ can be written as
@@
This is an off-shell symmetry with a corresponding Noether current. At large distances it interchanges the electric and magnetic parts of the Weyl tensor, and so is a Schwarzschild generalization of electric-magnetic duality. [^35] Note that this transformation keeps the background fixed, it does not map Schwarzschild into Taub–NUT. It is therefore directly a symmetry of linearized perturbations around Schwarzschild.
In the present context, this symmetry maps the worldline response operators (3.40) in the electric and magnetic sectors into each other, relating their coefficients. The astute reader, however, will note that the finite contributions in e.g., (3.87) and (3.118) differ. This would seem to be in conflict with the symmetry. However, we computed these quantities via dimensional regularization, and the symmetry (3.121) exists only in $D=4$. Since we have chosen a regularization scheme that breaks the symmetry, it is not surprising that the scheme-dependent parts of the final answers do not respect it. The existence of the symmetry implies the existence of a scheme that preserves the symmetry. Indeed, an important consistency check is that the scheme-independent parts of the response—the imaginary dissipative response and the coefficient of the logarithms—match exactly between the two sectors. In addition, we can certainly pick the scales $\mu$ in the two sectors differently so that the finite pieces agree, which would be a scheme that preserves the duality.
@@ 4 Conclusions
After matching to define the dynamical Love numbers of black holes, we find they have several interesting properties. They are nonzero, and have a characteristic running, where the coefficient of the logarithm in all cases is $\omega r_{s}$ times the corresponding dissipative response coefficient. One way to read this relation is that the beta function for the dynamical Love numbers is determined by the dissipative response. [^36] This is rather surprising because the long-distance running of the conservative response is sourced by the Schwarzschild potential and hence is universal, while the dissipative response is a finite-size effect particular to black holes. This relation appears to become more complicated at subleading order [74]. Another reading is that the discontinuity of the conservative part of $K$ is $\beta\omega/2$ times the dissipative part of $K$, where $\beta=4\pi r_{s}$ is the inverse black hole temperature. [^37] This looks suspiciously like a fluctuation–dissipation relation. However, it is a relation involving just the causal Green’s function. (Interestingly this relation continues to hold for the most transcendental parts of the response kernel at subleading orders in frequency, at least in the examples of [74].) It would be nice to understand these features of the logarithmic terms better. Another interesting feature is that, as expected from the symmetries of black hole perturbations, there exists a scheme in which the electric and magnetic dynamical Love numbers (along with the dissipative responses) are equal.
There are a number of future directions suggested by this study that it would be interesting to pursue. Perhaps the most obvious would be to extend the computation done here to general $\ell$ and (relatedly) to higher dimension. In principle there are only technical complications to doing so, and it would be very interesting to see if the patterns seen in the four-dimensional case survive there. Relatedly, one could compute dynamical responses to external electromagnetic fields, and consider the generalization to charged black hole responses. Perhaps the most phenomenologically interesting generalization would be to compute the full dynamical Love numbers for a spinning Kerr black hole. The main advances required are technical. In particular, resumming and regulating gravitational effects in the EFT is simplest in dimensional regularization, but gravitational couplings to spinning black holes in higher dimensions are complicated. At leading order, we can extract a Kerr black hole’s dynamical Love number by exploiting the fact that a Kerr black hole is a spinning Schwarzschild black hole. The coordinate transformation between the (rotating) black hole’s rest frame and the laboratory frame captures the leading-in-spin effect (similar to how a Kerr black hole’s zero frequency dissipation is related to the finite frequency Schwarzschild dissipation [46, 19]). The computation of the full response for Kerr is a problem that we hope to return to.
@@ Acknowledgements:
Thanks to John Joseph Carrasco, Lam Hui, Riccardo Penco, Chia-Hsien Shen, Mikhail Solon, and Zihan Zhou for helpful conversations, and to Giulia Isabella, Massimiliano M. Riva, and Jan Steinhoff for comments on the draft. AJ and DG are supported in part by DOE (HEP) Award DE-SC0025323. The work of MJR is supported through the NSF grant PHY-2309270. The research of LS has been funded by the French National Research Agency (ANR) under project ANR-24-CE31-1097-01. The work is supported by a PhD Joint Program of the International Research Center for Fundamental Scientific Discovery (IRC Discovery).
Note added:
Shortly after this paper, [128] appeared, which computes the odd-parity magnetic dynamical tidal response, with a slightly different perspective and formalism. Where our results overlap, they are in agreement.
@@ Appendix A Scalar dynamical response running
with $m$ a non-negative integer. [^38] In (A.8), the lower bound in the integral (which we chose to be $1$ for convenience) is completely immaterial, as this simply amounts to a redefinition of the integration constants in (2.21). Note that no $\log(x)/x^{\ell+1}$ scaling is contained in the homogeneous solutions of (2.17). From the properties of the hypergeometric function, each of these terms is of the form $\text{pol}_{1}(x)+\text{pol}_{2}(x)\times\log(x)$, at large $x$, where $\text{pol}_{1}(x)$ and $\text{pol}_{2}(x)$ are generic ordinary polynomials with positive powers of $x$. In conclusion, no $\log(x)/x^{\ell+1}$ term results from the $x\rightarrow\infty$ expansion of the $I_{n}$ integrals (A.6), for all $n$. The second integral in (A.4) can also be computed:
$\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)W_{\ell-1}(y)=-\sum_{k=0}^{2\ell-1}d_{k}\,q_{k}(x)-\sum_{k=1}^{2\ell-1}\frac{8}{k}d_{k}\,(x-1)^{k}+8\,c_{0}\log(x-1),$ (A.9)
where $q_{k}$ denotes the polynomial
$q_{k}(y)\equiv\left(\frac{12(y-1)}{1+k}+\frac{6(y-1)^{2}}{2+k}+\frac{(y-1)^{3}}{3+k}\right)(y-1)^{k}\,$, (A.10)
and the coefficients in the series expansion are
$d_{k}=\frac{1}{2^{k}}\sum_{n=0}^{k}\frac{\Gamma(\ell+1+n)\,\Gamma(\ell+1+k-n)}{\Gamma(\ell+1-n)\,\Gamma(\ell+1-(k-n))\,(n!)^{2}\,((k-n)!)^{2}}\bigl(\psi(\ell+1)-\psi(k-n+1)\bigr).$ (A.11)
From this expression, we can now make the following consideration. As noted before, for the first integral we argue that the same conclusion is trivially true for the second integral in (A.4), since both $P_{\ell}(y)$ and $W_{\ell-1}(y)$ are polynomials in (y1)(y-1). Therefore, there is no $\log(x)/x^{\ell+1}$ type of term resulting from the first integral in (2.21). Let us then focus on the second integral in (2.21):
@@
$\frac{B}{4}\int^{x}{\rm d}y\frac{(1+y)^{3}}{y-1}P_{\ell}(y)^{2}$ $=\frac{B}{4}\left(\sum_{m=0}^{2\ell}f_{m}\,q_{m}(x)+\sum_{m=1}^{2\ell}\frac{8}{m}f_{m}(x-1)^{m}+8\log(x-1)\right)$
$=2B\log(x-1)+\text{pol}(x),$ (A.13)
where the coefficients $f_{m}$ from the convolution of the series expansions of $P_{\ell}^{2}$ take the form
$f_{m}=\sum_{n=0}^{m}a_{n}\,a_{m-n}.$ (A.14)
@@ B.1 Far-zone scalar solution
$_{p}\tilde{F}_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right]\equiv\Big[\Gamma(b_{1})\cdots\Gamma(b_{q})\Big]^{-1}\,{}_{p}F_{q}\left[\begin{array}[]{c}a_{1}\,,\,\cdots\,,\,a_{p}\\[-2.0pt] b_{1}\,,\,\cdots\,,\,b_{q}\end{array}\bigg\rvert\,x\,\right].$ (B.17)
@@
To expand the Coulomb functions, we can then use their relations to spherical Bessel functions:
@@
Plugging this into the definition of $\phi^{(1)}_{\ell,\text{FZ}}=\frac{\omega}{z}(\xi^{(1)}_{\ell}+\frac{1}{2z}\xi^{(0)}_{\ell})$, we obtain the following $z\rightarrow 0$ behavior for $\phi_{\ell,\text{FZ}}^{(1)}$ [^43] We use $y_{0}^{\ell}=(2\ell-1)!!=\frac{(2\ell+1)!!}{2\ell+1}$ to simplify the expressions.:
@@
$\quad+c_{1}^{(1)}j_{0}^{\ell}z^{\ell}+O(z^{\ell+1})\Bigg]$ (B.27)
$=$ $\,\,\omega c_{2}^{(1)}\frac{(2\ell+1)!!}{2\ell+1}z^{-\ell-1}\left(1+\sum_{n=1}^{\ell-1}\frac{y_{n}^{\ell}}{y_{0}^{\ell}}z^{2n}\right)-\omega c_{1}^{(0)}\frac{\ell}{2(2\ell+1)!!}z^{\ell-1}+\frac{\omega c_{1}^{(1)}}{(2\ell+1)!!}z^{\ell}+O(z^{\ell+1}).$
@@ C.1 General relativistic solutions
$\Psi_{\text{Z,IZ}}^{\ell=3}(r)\underset{\frac{r}{r_{s}}\to\infty}{=}\frac{C}{50}\bigg(300\frac{r^{4}}{r_{s}^{4}}-190\frac{r^{3}}{r_{s}^{3}}-93\frac{r^{2}}{r_{s}^{2}}+\frac{279}{10}\frac{r}{r_{s}}+\frac{663}{100}-\frac{1\,989}{1\,000}\frac{r_{s}}{r}+\frac{5\,967}{10\,000}\frac{r_{s}^{2}}{r^{2}}-\frac{17\,901}{100\,000}\frac{r_{s}^{3}}{r^{3}}\bigg)$
@@
$\penalty 10000\ \hskip 2.84544pt+C\omega^{2}r_{s}^{2}\Bigg[-\frac{1}{3}\frac{r^{6}}{r_{s}^{6}}-\frac{11}{10}\frac{r^{5}}{r_{s}^{5}}-\frac{r^{4}}{r_{s}^{4}}\!\left(\pi^{2}+\frac{51\,917}{2\,100}-\frac{13}{7}\log\frac{r_{s}}{r}\right)+\frac{r^{3}}{r_{s}^{3}}\!\left(\frac{19\pi^{2}}{30}+\frac{399\,451}{21\,000}-\frac{247}{210}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 52.63777pt+\frac{r^{2}}{r_{s}^{2}}\!\left(\frac{31\pi^{2}}{100}+\frac{1\,634\,147}{210\,000}-\frac{403}{700}\log\frac{r_{s}}{r}\right)-\frac{r}{10^{3}r_{s}}\!\left(93\pi^{2}+\frac{2\,039\,147}{700}-\frac{1\,209}{7}\log\frac{r_{s}}{r}\right)$
@@
$\penalty 10000\ \hskip 52.63777pt+\frac{r_{s}^{2}}{10^{6}r^{2}}\!\left(-1\,989\pi^{2}+\frac{220\,495\,907}{2\,100}+\frac{25\,857}{7}\log\frac{r_{s}}{r}\right)$
$\penalty 10000\ \hskip 52.63777pt+\frac{r_{s}^{3}}{10^{7}r^{3}}\!\left(5\,967\pi^{2}+\frac{58\,238\,757\,859}{44\,100}-\frac{5\,232\,713}{21}\log\frac{r_{s}}{r}\right)\Bigg]+O\!\left(\frac{r_{s}^{4}}{r^{4}}\right),$ (C.3)
@@
$\Psi^{\ell=4}_{\text{Z,IZ}}(r)\underset{\frac{r}{r_{s}}\to\infty}{=}C\bigg(28\frac{r^{5}}{r_{s}^{5}}-\frac{119}{3}\frac{r^{4}}{r_{s}^{4}}+\frac{179}{18}\frac{r^{3}}{r_{s}^{3}}+\frac{361}{108}\frac{r^{2}}{r_{s}^{2}}-\frac{361}{648}\frac{r}{r_{s}}$
$\penalty 10000\ \hskip 2.84544pt+Ci\omega r_{s}\Bigg[\frac{1\,246}{15}\frac{r^{5}}{r_{s}^{5}}-\frac{10\,591}{90}\frac{r^{4}}{r_{s}^{4}}+\frac{15\,931}{540}\frac{r^{3}}{r_{s}^{3}}+\frac{32\,129}{3\,240}\frac{r^{2}}{r_{s}^{2}}-\frac{32\,129}{19\,440}\frac{r}{r_{s}}$
@@
$\penalty 10000\ \hskip 2.84544pt+C\omega^{2}r_{s}^{2}\Bigg[-\frac{14}{11}\frac{r^{7}}{r_{s}^{7}}-\frac{931}{330}\frac{r^{6}}{r_{s}^{6}}-\frac{r^{5}}{r_{s}^{5}}\!\left(\frac{14\pi^{2}}{3}+\frac{2\,798\,434}{17\,325}-\frac{3\,142}{495}\log\frac{r_{s}}{r}\right)$